OpenStudy (anonymous):

The probability that an event will happen is . What is the probability that the event will not happen? 0 3/8 5/8

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@LyricalDevonne

OpenStudy (anonymous):

@eliasperez191

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

please limit to one question per post

OpenStudy (anonymous):

@ganeshie8 OK i'll be sure to do that next time. I'm sorry. Are you still willing to help me?

ganeshie8 (ganeshie8):

which question ?

OpenStudy (anonymous):

The first one. I don't know to find the probability that an event won't happen.

ganeshie8 (ganeshie8):

I think there was something missing in the question.. looks there was some probolem with pasting..

ganeshie8 (ganeshie8):

The probability that an event will happen is ` `. What is the probability that the event will not happen?

ganeshie8 (ganeshie8):

i feel something is missing at that blank ` `

OpenStudy (anonymous):

It's 5/8. Sorry @ganeshie8

ganeshie8 (ganeshie8):

Okay, use below : P(not A) = 1 - P(A)

ganeshie8 (ganeshie8):

`probability that an event does not happen`= 1 - `probability that the happen happens`

OpenStudy (anonymous):

Am I able to figure out the equation off of that? I'm still confused.

ganeshie8 (ganeshie8):

you're given P(A) = 5/8

ganeshie8 (ganeshie8):

simply subtract it from 1 fro P(not A)

ganeshie8 (ganeshie8):

P(not A) = 1 - 5/8

OpenStudy (anonymous):

So 5/8-1=3/8

ganeshie8 (ganeshie8):

Yep!

OpenStudy (anonymous):

How come we subtract from one instead of any other number @ganeshie8

ganeshie8 (ganeshie8):

thats a very good question!

ganeshie8 (ganeshie8):

probability is a number between 0 and 1

ganeshie8 (ganeshie8):

and all the probabilities of outcomes add up to 1

ganeshie8 (ganeshie8):

since an even can either "happen" or "not happen", the sum of these probabilities will be 1 : P(A) + P(not A) = 1

OpenStudy (anonymous):

OK I feel like I have a better grasp on probability than I did before. Thanks! :)

ganeshie8 (ganeshie8):

no, i think im confusing u :) let me give u an example you will always win if you bet something like this : today it either rains or it doesnt rain

ganeshie8 (ganeshie8):

are there any other possibilities for it ? it either rains, or it doesnt rain. so P(rains today) + P(does not rain today) = 1